# Continuously tunable uniaxial strain control of van der Waals heterostructure devices

### v1

@@ Abstract
Uniaxial strain has been widely used as a powerful tool for investigating and controlling the properties of quantum materials. However, existing strain techniques have so far mostly been limited to use with bulk crystals. Although recent progress has been made in extending the application of strain to two-dimensional van der Waals (vdW) heterostructures, these techniques have been limited to optical characterization and extremely simple electrical device geometries. Here, we report a piezoelectric-based in-situ uniaxial strain technique enabling simultaneous electrical transport and optical spectroscopy characterization of dual-gated vdW heterostructure devices. Critically, our technique remains compatible with vdW heterostructure devices of arbitrary complexity fabricated on conventional silicon/silicon dioxide wafer substrates. We demonstrate a large and continuously tunable strain of up to $0.15$% at millikelvin temperatures, with larger strain values also likely achievable. We quantify the strain transmission from the silicon wafer to the vdW heterostructure, and further demonstrate the ability of strain to modify the electronic properties of twisted bilayer graphene. Our technique provides a highly versatile new method for exploring the effect of uniaxial strain on both the electrical and optical properties of vdW heterostructures, and can be easily extended to include additional characterization techniques.
@@ I Introduction
Strain is a powerful tool for directly manipulating the crystal lattice of materials, and consequently for tuning their electronic properties. For example, uniaxial strain can break the in-plane rotational symmetry of a lattice, potentially generating new electronic phases. Over the past decade, a powerful technique has been developed for applying continuously tunable uniaxial strain to three-dimensional bulk quantum materials using a device based on three piezoelectric stacks.Hicks et al. (2014a) This technique has been widely used to study and tune superconductivity,Hicks et al. (2014b); Steppke et al. (2017); Malinowski et al. (2020); Qian et al. (2021) topological phases,Mutch et al. (2019); Jo et al. (2019) and nematicityBartlett et al. (2021); Sanchez et al. (2021); Ghini et al. (2021); Wiecki et al. (2021) in a variety of bulk quantum materials, and to investigate their thermodynamic properties.Li et al. (2020); Ikeda et al. (2019) Over the past few years, van der Waals (vdW) heterostructures have emerged as exciting new platforms for exploring related strongly correlated and topological states in two dimensions. Flat electronic bands arise when two or more vdW sheets are properly stacked, and can host a wealth of emergent states including superconductivity, nematicity, ferromagnetism, generalized Wigner crystals, and both integer and fractional Chern insulators.Cao et al. (2018a); Yankowitz et al. (2019a); Chen et al. (2021); Cao et al. (2021); Fei et al. (2018); Serlin et al. (2020); Park et al. (2023) However, technical constraints have so far severely limited the extension of existing strain-tuning techniques to the study of these states in vdW heterostructures.
@@
Although there have been a number of approaches developed for applying strain to vdW materials, so far none are compatible with standard electrical transport measurements of dual-gated vdW devices assembled on a silicon/silicon dioxide (Si/SiO2) wafer substrate.Bunch et al. (2007); Conley et al. (2013); Lloyd et al. (2016); Deng, Sumant, and Berry (2018); Hou et al. (2019); Dai, Liu, and Zhang (2019); Kim et al. (2023) In this manuscript, we introduce a novel method that integrates traditional vdW heterostructure device fabrication on silicon substrates with a three-piezo-stack-based strain cell.Dean et al. (2010); Kim et al. (2016); Yankowitz et al. (2019b); Hicks et al. (2014a); Park et al. (2020); Bartlett et al. (2021) We quantify the induced strain in the active layer of the vdW heterostructure by Raman spectroscopy, and corroborate this value with estimates from metallic strain gauges either glued onto the piezo stack or evaporated onto the silicon wafer nearby the vdW device. We further characterize the strain transmission through multilayer vdW heterostructures by characterizing the strain induced hexagonal boron nitride (hBN) for different flake thicknesses and for twisted hBN multilayers. Our measurements indicate that strain is effectively transmitted from the silicon wafer into many layers of vdW flakes residing on top. Finally, we also showcase the capabilities of our strain technique by measuring the electrical transport properties of a twisted bilayer graphene (tBLG) device near the magic angle, and show that strain can induce large changes in the device resistivity. Our technique is compatible with additional experimental probes beyond optics and transport, and opens up new avenues for future experiments in vdW heterostructure devices.
@@ II Experimental methods
Our experimental design combines the well-established piezoelectric-based strain technique for bulk samplesHicks et al. (2014a); Park et al. (2020); Bartlett et al. (2021) with the standard dry-transfer technique for fabricating vdW heterostructure devices.Kim et al. (2016) We first discuss the challenges inherent to integrating these two techniques, and overview our solution of creating a bowtie-shape silicon substrate.
@@ II.1 Current challenges for straining 2D vdW devices
vdW heterostructure devices are typically fabricated on 500 $\mu$m thick Si/SiO2 wafer substrates.Kim et al. (2016); Cao et al. (2018b); Wang et al. (2013) A simple way to induce strain to the vdW heterostructure is to directly strain the substrate wafer. However, single crystalline silicon is a very brittle and rigid material, with a Young’s modulus of $\approx 130-180$ GPa,Hopcroft, Nix, and Kenny (2010) which severely limits the amount of strain that can be applied before the wafer shatters. Therefore, the challenge is to identify a suitable substrate that is flexible enough to induce high level of the strain, yet rigid enough to be compatible with the fabrication of state-of-the-art vdW heterostructure devices. One approach is to replace the silicon wafer with a flexible substrate, such as a thin metal or a polymer. Wang et al. (2019); Mohiuddin et al. (2009); Wu et al. (2014) For example, large tensile strains were previously induced in a vdW material by using a three-point bending geometry with a polyimide-coated phosphor bronze substrate.Wang et al. (2019) However, a continuous tuning of strain was not demonstrated, and it is not clear how this setup can induce a compressive strain.
@@
However, the very high Young’s modulus of silicon still presents a significant challenge. Without any modification, inducing a one percent strain in a regular silicon wafer requires a force that far exceeds the blocking force of the typical piezostacks used in commercial strain cells. Several approaches have been developed so far in attempt to circumvent this issue. One approach is to suspend a sample across a micron-meter size gap ($\sim$3-5 $\mu$m), which is created by carefully cleaving the silicon wafer.Cenker et al. (2022) Although this approach can result in the application of stain exceeding 1% to a thin vdW flake, it is extraordinarily difficult to fabricate complex vdW device architectures that are suspend across the gap.
@@
Another approach is to transfer the vdW onto a thin silicon wafer strip, ideally only a few hundred micrometers wide. The small cross-section of the thin strip greatly reduces the amount of force required to strain the substrate. Cenker et al. (2022); Hwangbo et al. (2023) However, the extremely small size of thin strip makes the fabrication of conventional vdW heterostructure devices very difficult. To address this issue, we developed a new approach in which we first shape the silicon wafer into a bowtie shape. This geometry is inspired by a similar titanium platform structure used in Ref. Park et al., 2020; Bartlett et al., 2021 to strain bulk crystals. The key advantage of this design is the narrow bridge of the bowtie greatly reduces the amount of force required to induce strain in that region, while the large area at two ends of the wafer simultaneously enables the fabrication of arbitrarily complex vdW heterostructure devices with top and bottom gates and many electrical contacts.
@@
Figure 1: (a) Schematic of the strain apparatus assembly. The light gray frame is constructed of titanium. The piezo stacks (dark gray) have dimensions of $5\times 5\times 9$ mm.pie A PCB is affixed onto the side of the strain cell on the titanium frame, and hovers above the nearby piezostack. A second PCB is affixed onto the other side of the strain cell, but is not shown here for clarity. The PCB is used to connect electrical wires to both the sample and the wiring of the cryostat insert. The strain cell is affixed to a PCB carrier by a beam holder, and the entire assembly can be mounted onto various cryostat inserts. (b) Top view of the apparatus. The chip carrier (orange) is tightly affixed onto the titanium frame by two screws. The bowtie-shaped 50 $\mu$m thick silicon chip (purple) is epoxied onto the chip carrier with Stycast 2850FT epoxy (black). The top surface is partially enveloped by the epoxy, denoted by the white dashed lines. The 25 $\mu$m gold wires (yellow) are bonded to the PCB pads (green) using silver paste. A commercial strain gauge is glued on the central piezo stack. (c) Side-view photograph of the strain cell attached to a Bluefors dilution refrigerator top-loading insert. (d) Top-view photograph of the strain cell. A temperature sensor made of carbon composition resistor Samkharadze, Kumar, and Csáthy (2010) was adhered to the side wall of the titanium block with GE vanish to monitor the local temperature of the strain cell in the cryostat.
@@ II.2 The design of strain cell for vdW 2D device
Figure 2: (a) Cartoon schematic illustrating the laser cutting process used to create a 50 $\mu$m thick bowtie-shaped silicon chip. The thin silicon wafer (gray) is first glued onto a thicker silicon substrate wafer with PMMA. The 285 nm thick SiO2 capping layer faces downward to avoid overheating during laser cutting. By carefully adjusting the power, the laser can cut throughout the Si/SiO2 layers without degrading the SiO2 surface. The dashed lines indicate the laser cutting path. The light blue indicates regions of the silicon wafer are removed after cutting. The outer square frame is used to isolate many bowtie-shaped chips from a single large-area wafer. (b) Schematic of a standard dry transfer process of a vdW heterostructure onto the clean SiO2 layer, located at the center bridge of the pre-cut silicon chip, as shown in (a). (c) Layout of the final silicon chip after electron beam lithography, plasma etching, and metal deposition (left). Optical image of the central bridge region (middle), and the vdW heterostructure device with a meandering gold pattern strain gauge (right). The central bridge width is about 0.3 mm. The white scale bars represent 100 $\mu$m (middle) and 25 $\mu$m (right), respectively. (d) Diagram of the silicon chip glued on top of the chip carrier, which is pre-fixed by two screws on the strain cell. The epoxied area regions are outlined by the white dashed lines. The two outside arms (marked by the black crosses) are cut after the epoxy has cured.
@@
We use a home-built piezoelectric strain cell with a design similar to the one introduced by Hicks et al.Hicks et al. (2014a) As depicted in Fig. 1, the device comprises three piezo stacks ($5\times 5\times 9$ mm, with high stiffness constant and blocking forcepie ) and custom-machined titanium frames. We further include a beam holder that affixes the strain cell to a printed circuit board (PCB) sample mount, which allows us to easily wire up the strain cell for use in different cryostats (we have so far used a PPMS from Quantum Design, Inc., and a top-loading dilution refrigerator from Blueforsthe ). Viewed from above the strain cell, the bowtie-shaped Si chip (colored purple) is glued with Stycast epoxysty onto a titanium chip carrier (colored orange), which is prefixed onto the titanium sample mount by screws on each end. The glue covers roughly half of the top of the tab area of the bowtie to ensure robust mechanical anchoring, while leaving enough space for around eight electrical contacts on each side. The gold electrical pads are wired to a custom PCB that is mounted on the side of the strain cell. A commercial strain gaugestr is glued onto one of the piezo stacks, and is used to estimate the displacement of the titanium chip carrier by the piezo stacks. A small carbon resistor is adhered to the titanium frame of the strain cell in order to monitor the local temperatureSamkharadze, Kumar, and Csáthy (2010).
@@
We fabricate our vdW heterostructure devices atop a 50 $\mu$m thick silicon wafer with a 285 nm SiO2 capping layer.IWS This wafer is ten times thinner than those conventionally used for vdW heterostructure devices. The reduced thickness is important for lowering the spring constant and enabling the application of larger strain.Cenker et al. (2022) We first laser-cut two rectangular holes out of the wafer in order to create the bowtie shape. In contrast to previous methods used for laser cutting titanium and quartz platesPark et al. (2020); Bartlett et al. (2021), additional steps are needed to protect the surface of SiO2 layer when laser cutting the silicon wafer since the vdW heterostructure is extremely sensitive to the surface cleanliness. Therefore, prior to the cutting, we glue the thin wafer atop a thicker Si substrate using polymethyl methacrylate (PMMA) as the bonding adhesive. The SiO2 surface of the thin wafer is arranged to face downward, in contact with the PMMA adhesive. There are two benefits of doing this: first, to provide mechanical stability for the fragile thin wafer, and second, to protect the SiO2 layer from silicon dust generated during the laser cutting process. The substrate wafer is affixed onto an aluminum plate with double-sided tape, and the plate is mounted on the laser cutting board. We employed a laser cutting system (LPKF ProtoLaser) to cut the silicon wafer. We note that the laser power is critical since over cutting can also damage the SiO2 layer, or result in poor detachment from the PMMA. Here, we use a laser power of 1.2 watts with a beam diameter of 10 $\mu$m and 1000 times cut repetition for each pattern. The square silicon wafer is cut into the designated pattern outlined by the dashed lines shown in Fig. 2(a). Note that this pattern includes two outer bridges in addition to the central bridge of the bow-tie, as shown in Fig. 2(b). These outer bridges are temporary, but they are essential for providing mechanical stability during the transfer and fabrication of vdW heterostructure device. These two outer bridges are cut by hand after mounting the wafer onto the piezoelectric strain cell. The bridges of the bowtie are aligned along <110> directions of a (100) wafer.
@@
After laser cutting, the small three-bridge-patterned thin silicon chips are detached from the Si substrate wafer by sonicating in acetone. The three-bridge-patterned chip is then glued to another rigid silicon substrate with PMMA, this time with its SiO2 layer facing upwards (Fig. 2(b)). A vdW heterostructure is assembled layer-by-layer using the standard dry-transfer technique with a poly-carbonate (PC) stamp. After the stack is assembled, the PC stamp is melted onto the central bridge of the thin silicon wafer and soaked in chloroform dissolve the PC, leaving behind only the vdW heterostructure. The vdW heterostructure is then processed into a Hall bar geometry and contacted electrically following a standard series of steps involving electron beam lithography, plasma etching, and metal deposition. We arrange the Hall bar such that it is aligned along the direction of uniaxial strain. Figure 2(c) shows a typical three-bridge chip with a tBLG device at its center. In addition to the gold electrodes connected to the device, we also evaporate gold along the laser-cut edges of the chip. These gold strips are essential, otherwise the rough wafer edges resulting from laser cutting can prevent a smooth metal liftoff procedure. Additionally, the roughness can make the wafer more prone to breaking under tensile strain, but these can be smoothed and potentially strengthened through chemical etching. We also evaporate a meander-shaped gold wire nearby the device, which serves as a custom local strain gauge. The three-bridge silicon chip is then epoxied to the titanium chip carrier pre-mounted on the strain cell, as described above. The final step is to cut the outer two bridges (at positions indicated by the black crosses in Fig. 2(d)) using a scalpel blade after the epoxy has cured.

### v2

@@ Abstract
Uniaxial strain has been widely used as a powerful tool for investigating and controlling the properties of quantum materials. However, existing strain techniques have so far mostly been limited to use with bulk crystals. Although recent progress has been made in extending the application of strain to two-dimensional van der Waals (vdW) heterostructures, these techniques have been limited to optical characterization and extremely simple electrical device geometries. Here, we report a piezoelectric-based in situ uniaxial strain technique enabling simultaneous electrical transport and optical spectroscopy characterization of dual-gated vdW heterostructure devices. Critically, our technique remains compatible with vdW heterostructure devices of arbitrary complexity fabricated on conventional silicon/silicon dioxide wafer substrates. We demonstrate a large and continuously tunable strain of up to $-0.15\%$ at millikelvin temperatures, with larger strain values also likely achievable. We quantify the strain transmission from the silicon wafer to the vdW heterostructure, and further demonstrate the ability of strain to modify the electronic properties of twisted bilayer graphene. Our technique provides a highly versatile new method for exploring the effect of uniaxial strain on both the electrical and optical properties of vdW heterostructures, and can be easily extended to include additional characterization techniques.
@@ I Introduction
Strain is a powerful tool for directly manipulating the crystal lattice of materials, and consequently for tuning their electronic properties. For example, uniaxial strain can break the in-plane rotational symmetry of a lattice, potentially generating new electronic phases. Over the past decade, a powerful technique has been developed for applying continuously tunable uniaxial strain to three-dimensional bulk quantum materials using a device based on three piezoelectric stacks.Hicks et al. (2014a) This technique has been widely used to study and tune superconductivity,Hicks et al. (2014b); Steppke et al. (2017); Malinowski et al. (2020); Qian et al. (2021) topological phases,Mutch et al. (2019); Jo et al. (2019) and nematicityBartlett et al. (2021); Sanchez et al. (2021); Ghini et al. (2021); Wiecki et al. (2021) in a variety of bulk quantum materials and to investigate their thermodynamic properties.Li et al. (2020); Ikeda et al. (2019) Over the past few years, van der Waals (vdW) heterostructures have emerged as exciting new platforms for exploring related strongly correlated and topological states in two dimensions. Flat electronic bands arise when two or more vdW sheets are properly stacked and can host a wealth of emergent states including superconductivity, nematicity, ferromagnetism, generalized Wigner crystals, and both integer and fractional Chern insulators.Cao et al. (2018a); Yankowitz et al. (2019a); Chen et al. (2021); Cao et al. (2021); Fei et al. (2018); Serlin et al. (2020); Balents et al. (2020); Andrei and MacDonald (2020); Park et al. (2023); Cao et al. (2020); Polshyn et al. (2019); Jaoui et al. (2022) However, technical constraints have so far severely limited the extension of existing strain-tuning techniques to the study of these states in vdW heterostructures.
@@
Although there have been a number of approaches developed for applying strain to vdW materials, so far none are compatible with standard electrical transport measurements of dual-gated vdW devices assembled on a silicon/silicon dioxide (Si/SiO2) wafer substrate.Bunch et al. (2007); Conley et al. (2013); Lloyd et al. (2016); Deng, Sumant, and Berry (2018); Hou et al. (2019); Dai, Liu, and Zhang (2019); Kim et al. (2023) Prior methods have mostly focused either on suspended vdW samples, which are mechanically fragile, or devices with limited numbers of gate and contact electrodes. In this paper, we introduce a novel method that integrates traditional vdW heterostructure device fabrication on silicon substrates with a three-piezo-stack-based strain cell.Dean et al. (2010); Kim et al. (2016); Yankowitz et al. (2019b); Hicks et al. (2014a); Park et al. (2020); Bartlett et al. (2021) We quantify the induced strain in the active layer of the vdW heterostructure by Raman spectroscopy and corroborate this value with estimates from metallic strain gauges either glued onto the piezo stack or evaporated onto the silicon wafer near the vdW device. We further characterize the strain transmission through multilayer vdW heterostructures by characterizing the strain induced hexagonal boron nitride (hBN) for different flake thicknesses and for twisted hBN multilayers. Our measurements indicate that strain is effectively transmitted from the silicon wafer into many layers of vdW flakes residing on top. Finally, we also showcase the capabilities of our strain technique by measuring the electrical transport properties of a twisted bilayer graphene (tBLG) device near the magic angle and show that strain can induce large changes in the device resistivity. Our technique is compatible with additional experimental probes beyond optics and transport and opens up new avenues for future experiments in vdW heterostructure devices.
@@ II Experimental methods
Our experimental design combines the well-established piezoelectric-based strain technique for bulk samplesHicks et al. (2014a); Park et al. (2020); Bartlett et al. (2021) with the standard dry-transfer technique for fabricating vdW heterostructure devices.Kim et al. (2016) We first discuss the challenges inherent to integrating these two techniques, and overview our solution of creating a bowtie-shaped silicon substrate.
@@ II.1 Current challenges for straining 2D vdW devices
vdW heterostructure devices are typically fabricated on 500 $\mu$m thick Si/SiO2 wafer substrates.Kim et al. (2016); Cao et al. (2018b); Wang et al. (2013) A simple way to induce strain to the vdW heterostructure is to directly strain the substrate wafer. However, single crystalline silicon is a rigid material, with Young’s modulus of $\approx 130-180$ GPa.Hopcroft, Nix, and Kenny (2010) Hence, the force required to strain a millimeter size wafer to $\approx 1$% is approximately 750 N, far exceeds what can be achieved in existing piezo-based strain devices. Therefore, it remains an open challenge to identify a suitable substrate that is flexible enough to achieve high levels of the strain, yet rigid enough to be compatible with the fabrication of state-of-the-art vdW heterostructure devices. One approach is to replace the silicon wafer with a flexible substrate, such as a thin metal or a polymer. Wang et al. (2019); Mohiuddin et al. (2009); Wu et al. (2014); Zhang et al. (2017); Ma et al. (2020) For example, large tensile and compressive strains were previously induced in a vdW material by using a two- or three-point bending geometry with flexible substrates such as polyimide-coated phosphor bronze or soft polyethylene-based thin films. Wang et al. (2019); Zhang et al. (2017); Ma et al. (2020) However, bending the substrate can additionally induce an unintentional vertical strain gradient in multilayer vdW heterostructures.
@@
However, the very high Young’s modulus of silicon still presents a significant challenge. Without any modification, inducing a one percent strain in a regular silicon wafer requires a force that far exceeds the blocking force of the typical piezo stacks used in commercial strain cells. Several approaches have been developed so far in an attempt to circumvent this issue. One approach is to suspend a sample across a micrometer size gap ($\sim$3-5 $\mu$m), which is created by carefully cleaving the silicon wafer.Cenker et al. (2022) Although this approach can result in the application of strain exceeding 1% to a thin vdW flake, it is extraordinarily difficult to fabricate complex vdW device architectures that are suspended across the gap.
@@
Another approach is to transfer the vdW onto a thin silicon wafer strip, ideally only a few hundred micrometers wide. The small cross section of the thin strip greatly reduces the amount of force required to strain the substrate. Cenker et al. (2022); Hwangbo et al. (2023) However, the extremely small size of thin strip makes the fabrication of conventional vdW heterostructure devices very difficult. To address this issue, we developed a new approach in which we first shape the silicon wafer into a bowtie shape. This geometry is inspired by a similar titanium platform structure used in Refs. Bartlett et al., 2021 and Park et al., 2020 to strain bulk crystals. The key advantage of this design is the narrow bridge of the bowtie greatly reduces the amount of force required to induce strain in that region, while the large area at two ends of the wafer simultaneously enables the fabrication of arbitrarily complex vdW heterostructure devices with top and bottom gates and many electrical contacts.
@@
Figure 1: (a) Schematic of the strain apparatus assembly. The light gray frame is constructed of titanium. The piezo stacks (dark gray) have dimensions of $5\times 5\times 9$ mm.pie A PCB is affixed onto the side of the strain cell on the titanium frame and hovers above the nearby piezo stack. A second PCB is affixed onto the other side of the strain cell, but is not shown here for clarity. The PCB is used to connect electrical wires to both the sample and the wiring of the cryostat insert. The strain cell is affixed to a PCB carrier by a beam holder, and the entire assembly can be mounted onto various cryostat inserts. (b) Top view of the apparatus. The chip carrier (orange) is tightly affixed onto the titanium frame by two screws. The bowtie-shaped 50 $\mu$m thick silicon chip (purple) is epoxied onto the chip carrier with Stycast 2850FT epoxy (black). The top surface is partially enveloped by the epoxy, denoted by the white dashed lines. The 25 $\mu$m gold wires (yellow) are bonded to the PCB pads (green) using silver paste. A commercial strain gauge is glued on the central piezo stack. (c) Side-view photograph of the strain cell attached to a Bluefors dilution refrigerator top-loading insert. (d) Top-view photograph of the strain cell. A temperature sensor made of carbon composition resistor Samkharadze, Kumar, and Csáthy (2010) was adhered to the sidewall of the titanium block with GE vanish to monitor the local temperature of the strain cell in the cryostat.
@@ II.2 The design of strain cell for vdW 2D device
Figure 2: (a) Cartoon schematic illustrating the laser cutting process used to create a 50 $\mu$m thick bowtie-shaped silicon chip. The thin silicon wafer (gray) is first glued onto a thicker silicon substrate wafer with PMMA. The 285 nm thick SiO2 capping layer faces downward to avoid overheating during laser cutting. By carefully adjusting the power, the laser can cut throughout the Si/SiO2 layers without degrading the SiO2 surface. The dashed lines indicate the laser cutting path. The light blue indicates that regions of the silicon wafer are removed after cutting. The outer square frame is used to isolate many bowtie-shaped chips from a single large-area wafer. (b) Schematic of a standard dry transfer process of a vdW heterostructure onto the clean SiO2 layer, located at the center bridge of the pre-cut silicon chip, as shown in (a). (c) Layout of the final silicon chip after electron beam lithography, plasma etching, and metal deposition (left). Optical image of the central bridge region (middle), and the vdW heterostructure device with a meandering gold pattern strain gauge (right). The central bridge width is about 0.3 mm. The white scale bars represent 100 $\mu$m (middle) and 25 $\mu$m (right), respectively. (d) Diagram of the silicon chip glued on top of the chip carrier, which is pre-fixed by two screws on the strain cell. The epoxied area regions are outlined by the white dashed lines. The two outside arms (marked by the black crosses) are cut after the epoxy has cured.
@@
We use a home-built piezoelectric strain cell with a design similar to the one introduced by Hicks et al.Hicks et al. (2014a) As depicted in Fig. 1, the device comprises three piezo stacks ($5\times 5\times 9$ mm, with high stiffness constant and blocking forcepie ) and custom-machined titanium frames. We further include a beam holder that affixes the strain cell to a printed circuit board (PCB) sample mount, which allows us to easily wire up the strain cell for use in different cryostats (we have so far used a PPMS from Quantum Design, Inc., and a top-loading dilution refrigerator from Blueforsthe ). Viewed from above the strain cell, the bowtie-shaped silicon chip (colored purple) is glued with Stycast epoxysty onto a titanium chip carrier (colored orange), which is prefixed onto the titanium sample mount by screws on each end. The glue covers roughly half of the top of the tab area of the bowtie to ensure robust mechanical anchoring, while leaving enough space for around eight electrical contacts on each side. The gold electrical pads are wired to a custom PCB that is mounted on the side of the strain cell. A commercial strain gaugestr is glued onto one of the piezo stacks and is used to estimate the displacement of the titanium chip carrier by the piezo stacks. A small carbon resistor is adhered to the titanium frame of the strain cell in order to monitor the local temperatureSamkharadze, Kumar, and Csáthy (2010).
@@
We fabricate our vdW heterostructure devices atop a 50 $\mu$m thick silicon wafer with a 285 nm SiO2 capping layer.IWS This wafer is ten times thinner than those conventionally used for vdW heterostructure devices. The reduced thickness is important for lowering the spring constant and enabling the application of larger strain.Cenker et al. (2022) We first laser-cut two rectangular holes out of the wafer in order to create the bowtie shape. In contrast to previous methods used for laser cutting titanium and quartz platesPark et al. (2020); Bartlett et al. (2021), additional steps are needed to protect the surface of SiO2 layer when laser cutting the silicon wafer since the vdW heterostructure is extremely sensitive to the surface cleanliness. Therefore, prior to the cutting, we glue the thin wafer atop a thicker silicon substrate using polymethyl methacrylate (PMMA) as the bonding adhesive. The SiO2 surface of the thin wafer is arranged to face downward, in contact with the PMMA adhesive. There are two benefits of doing this: first, to provide mechanical stability for the fragile thin wafer, and second, to protect the SiO2 layer from silicon dust generated during the laser cutting process. The substrate wafer is affixed onto an aluminum plate with double-sided tape, and the plate is mounted on the laser cutting board. We employed a laser cutting system (LPKF ProtoLaser) to cut the silicon wafer. We note that the laser power is critical since over cutting can also damage the SiO2 layer or result in poor detachment from the PMMA. Here, we use a laser power of 1.2 watts with a beam diameter of 10 $\mu$m and 1000 times cut repetition for each pattern. The square silicon wafer is cut into the designated pattern outlined by the dashed lines shown in Fig. 2(a). Note that this pattern includes two outer bridges in addition to the central bridge of the bowtie, as shown in Fig. 2(b). These outer bridges are temporary, but they are essential for providing mechanical stability during the transfer and fabrication of vdW heterostructure device. These two outer bridges are cut by hand after mounting the wafer onto the piezoelectric strain cell. The bridges of the bowtie are aligned along <110> directions of a (100) wafer.
@@
After laser cutting, the small three-bridge-patterned thin silicon chips are detached from the Si substrate wafer by sonicating in acetone. The three-bridge-patterned chip is then glued to another rigid silicon substrate with PMMA, this time with its SiO2 layer facing upwards [Fig. 2(b)]. A vdW heterostructure is assembled layer-by-layer using the standard dry-transfer technique with a poly-carbonate (PC) stamp. After the stack is assembled, the PC stamp is melted onto the central bridge of the thin silicon wafer and soaked in chloroform to dissolve the PC, leaving behind only the vdW heterostructure. Since the patterned wafer detaches from the thicker Si/SiO2 substrate wafer in chloroform, it must be re-affixed to a new rigid substrate for further processing. The vdW heterostructure is then processed into a Hall bar geometry and contacted electrically following a standard series of steps involving electron beam lithography, plasma etching, and metal deposition. The bowtie pattern created in the silicon wafer does not impede PMMA spin coating and electron-beam lithography; thus, these procedures are essentially the same as in the standard fabrication process for vdW heterostructures. We arrange the Hall bar such that it is aligned along the direction of uniaxial strain. Figure 2(c) shows a typical three-bridge chip with a tBLG device at its center. In addition to the gold electrodes connected to the device, we also evaporate gold along the laser-cut edges of the chip. These gold strips are essential; otherwise, the rough wafer edges resulting from laser cutting can prevent a smooth metal liftoff procedure. Additionally, the roughness can make the wafer more prone to breaking under tensile strain, but these can be smoothed and potentially strengthened through chemical etching. We also evaporate a meander-shaped gold wire near the device, which serves as a custom local strain gauge. The three-bridge silicon chip is then epoxied to the titanium chip carrier pre-mounted on the strain cell, as described above. The final step is to cut the outer two bridges [at positions indicated by the black crosses in Fig. 2(d)] using a scalpel blade after the epoxy has cured.